# Projection-free computation of robust controllable sets with constrained zonotopes

### v1

@@ 2.3 Robust controllable set
$\mathcal{K}=\left\{x_{0}\in\mathcal{X}_{0}\middle|\forall t\in\mathbb{N}_{[0:T-1]},\ \exists u_{t}\in\mathcal{U}_{t},\ \forall w_{t}\in\mathcal{W}_{t},\ x_{t+1}=A_{t}x_{t}+B_{t}u_{t}+F_{t}w_{t},\ x_{t}\in\mathcal{X}_{t},\ x_{T}\in\mathcal{G}\right\}.$
Informally, the $T$-step RC set $\mathcal{K}\subset\mathbb{R}^{n}$ is the set of initial states that, despite the additive disturbance $w_{t}\in\mathcal{W}_{t}$, can be steered using $u_{t}\in\mathcal{U}_{t}$ to reach the goal set $\mathcal{G}$ at time step $T$, while staying within the sequence of state constraints ${\{\mathcal{X}_{t}\}}_{t=0}^{T-1}$ at all intermediate time steps. The $T$-step RC set is also known in the literature as robust reachability of target tube [5], backward reachable set [9], or backward reach-avoid set [10]. We can compute $\mathcal{K}$ via a set recursion,
@@
$\operatorname{Pre}(\mathcal{K}_{t+1})$$\triangleq\{x:A_{t}x\in(\mathcal{K}_{t+1}\ominus F_{t}\mathcal{W})\oplus(-B_{t}\mathcal{U})\},$ (17b)
with $\mathcal{K}_{T}\triangleq\mathcal{G}$ and $\mathcal{K}=\mathcal{K}_{0}$. Observe that (2.3) uses all set operations in (2.2). For convex polyhedral/polytopic $\mathcal{X},\mathcal{G},\mathcal{S},\mathcal{U}$, and $\mathcal{W}$, we can compute RC sets using polyhedral computational geometry [4, 10, 3]. However, we typically face numerical challenges when implementing (2.3) using polytopes for high-dimensional systems (16), or over long horizons $T$, primarily due to projection required by the combination of Minkowski sum and intersection operations [3, 15].
On the other hand, constrained zonotopes provide a tractable, exact solution when $\mathcal{W}=\emptyset$ using (2.2[6, 7, 9]. Additionally, for a zonotopic $\mathcal{W}$[9] uses the two-stage approach reviewed in Section 2.2 to propose an inner-approximation of $\mathcal{K}$. In this paper, we use our solution to Problem 1 to inner-approximate $\mathcal{K}$ for a broader class of disturbance sets $\mathcal{W}$ that are symmetric, convex, and compact.
@@
Compute a constrained zonotopic $\mathcal{K}^{-}$ that inner-approximates $\mathcal{K}$ defined in Definition 2, where for every $t\in\mathbb{N}_{[0:T-1]}$, $\mathcal{U}_{t}$ and $\mathcal{G}$ are convex polytopes, $\mathcal{W}_{t}$ are symmetric, convex, and compact sets, and either 1) $\mathcal{X}_{t}$ are convex polyhedra and $A_{t}$ are invertible, or 2) $\mathcal{X}_{t}$ are convex polytopes. Additionally, characterize the representation complexity of $\mathcal{K}^{-}$ in both cases.
In both settings considered, the exact RC sets are known to be polytopes [3], and hence are representable via constrained zonotopes. We also use the proposed outer-approximation to the Pontryagin difference to outer-approximate $\mathcal{K}_{t}$ when using (2.3).
@@ 3 Inner-approximation of Pontryagin difference
Given a constrained zonotopic minuend $C\subset\mathbb{R}^{n}$ and a symmetric, convex, and compact subtrahend $\mathcal{S}\subset\mathbb{R}^{n}$, we address the inner-approximation part of Problem 1 by characterizing a constrained zonotope,
@@
$G_{M^{-}}=G_{C}D,\qquad c_{M^{-}}=c_{C}-c_{S},\qquad A_{M^{-}}=A_{C}D,\qquad b_{M^{-}}=b_{C},$ (19a)
for a specific $c_{S}\in\mathcal{S}$, similarly to (14). By construction, the representation complexity of $\mathcal{M}^{-}$ is $\mathscr{C}(\mathcal{M}^{-})=(n,N_{C},M_{C})=\mathscr{C}(\mathcal{C})$.
We now state a mild assumption on the minuend $\mathcal{C}$ that we will use to compute $\mathcal{M}^{-}$.
Subtrahend $\mathcal{S}$ Definition of the diagonal matrix $D=\operatorname{diag}_{i\in\mathbb{N}_{[1:N_{C}]}}(D_{ii})$ Result Convex and compact $\mathcal{S}$ that is symmetric about $c_{S}\in\mathbb{R}^{n}$ $D_{ii}=1-\rho_{\mathcal{S}_{0}}\left({e_{i}^{\top}{[G_{C};A_{C}]^{\dagger}[I_{n};0_{M_{C}\times n}]}}\right)$ for $\mathcal{S}_{0}\triangleq\mathcal{S}-c_{S}$ Thm. 2 Convex union of symmetric intervals (1c) $\mathcal{S}=\{G_{S}\xi+c_{S}:\|\xi\|_{1}\leq 1\}$ $D_{ii}=1-{\left\|{e_{i}^{\top}[G_{C};A_{C}]^{\dagger}[G_{S};0_{M_{C}\times N_{S}}]}\right\|}_{\infty}$  Zonotope (1d) $\mathcal{S}=\{G_{S}\xi+c_{S}:\|\xi\|_{\infty}\leq 1\}$ $D_{ii}=1-{\left\|{e_{i}^{\top}[G_{C};A_{C}]^{\dagger}[G_{S};0_{M_{C}\times N_{S}}]}\right\|}_{1}$ Corr. 1 Ellipsoid (1e) $\mathcal{S}=\{G_{S}\xi+c_{S}:\|\xi\|_{2}\leq 1\}$ $D_{ii}=1-{\left\|{e_{i}^{\top}[G_{C};A_{C}]^{\dagger}[G_{S};0_{M_{C}\times n}]}\right\|}_{2}$
@@
Assumption 1.
The minuend $\mathcal{C}$ is full-dimensional.
Observe that Assumption 1 on the minuend $\mathcal{C}$ is not restrictive, as illustrated with the following two observations. First, $\mathcal{C}\ominus\mathcal{S}=\emptyset$, whenever the affine dimension of $\mathcal{S}$ is strictly greater than the affine dimension of $\mathcal{C}$ by (6d), since $x+\mathcal{S}$ can not be contained in $\mathcal{C}$ for any $x\in\mathbb{R}^{n}$ in such a scenario. Second, consider the case where both minuend and subtrahend are not full-dimensional and there exists a matrix $T\in\mathbb{R}^{n\times n^{\prime}}$ with $n>n^{\prime}$ and $\text{rank}(T)=n^{\prime}$ such that $\mathcal{C}=T\mathcal{C}^{\prime}$ and $\mathcal{S}=T\mathcal{S}^{\prime}$ for any $\mathcal{C}^{\prime},\mathcal{S}^{\prime}\subset\mathbb{R}^{n^{\prime}}$. Then, by [23, Thm. 2.1.viii],
In such a case, it suffices to assume that $\mathcal{C}^{\prime}$ is full-dimensional in order to inner-approximate $\mathcal{C}\ominus\mathcal{S}$.
We first show that Assumption 1 enables an equivalent representation of the minuend. We then address the inner-approximation part of Problem 1, and conclude the section by relating our approach with the two-stage approach in [9] for zonotopic subtrahends. Table 1 summarizes the computation of $D$ for various types of sets considered in this paper.
3.1 MinRow representation for constrained zonotopes
We now show that Assumption 1 enables a useful, equivalent representation for a constrained zonotope (1b).
@@ 3.1 MinRow representation for constrained zonotopes
(MinRow Representation) A constrained zonotope $\mathcal{C}=(G_{C},c_{C},A_{C},b_{C})$ has a MinRow representation when $[G_{C};A_{C}]$ has full row rank.
A constrained zonotope (1b) in MinRow representation ensures that the matrix $[G_{C};A_{C}]$ has a well-defined right pseudoinverse ${[G_{C};A_{C}]}^{\dagger}$. However, the MinRow representation for a constrained zonotope is not unique, since one can preserve the rank of $[G_{C};A_{C}]$ without altering $\mathcal{C}$ by appropriately modifying $G_{C}$ and $A_{C}$ (e.g., increase $N_{C}$ by adding columns of zeros to $G_{C}$ and $A_{C}$).
@@
(Rank implications) Let Assumption 1 hold. Then, $G_{C}$ has full row rank, and $\mathcal{C}$ has a MinRow representation.
See Appendix A.1 for the proof of Proposition 1.
Computing a MinRow representation of a full-dimensional constrained zonotope $\mathcal{C}$: By Proposition 1, a full-dimensional $\mathcal{C}$ does not have a MinRow representation only if $A_{C}$ has linearly dependent rows. For such a set $\mathcal{C}$, $[A_{C},b_{C}]$ must also have linearly dependent rows since $\mathcal{C}$ is non-empty (see Appendix A.1 for more details). We compute a MinRow representation of $\mathcal{C}$ in the form of $\mathcal{C}=(G_{C},c_{C},A_{C}^{\prime},b_{C}^{\prime})$, where $[A_{C}^{\prime},b_{C}^{\prime}]$ is a set of linearly independent rows of the matrix $[A_{C},b_{C}]$ such that $\operatorname{rank}([A_{C}^{\prime},b_{C}^{\prime}])=\operatorname{rank}([A_{C},b_{C}])$. The rest of the paper assumes that a MinRow representation is used whenever a constrained zonotope satisfies Assumption 1.
@@ 3.2 Inner-approximation of the Pontryagin difference
We are now ready to address the inner-approximation part of Problem 1. First, we recall the following lemma.
Lemma 1.
[9, Lemma 2(ii)] For any set $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathbb{R}^{n}$ and $G\in\mathbb{R}^{m\times n}$ with $m<n$, $G(\mathcal{S}_{1}\ominus\mathcal{S}_{2})\subseteq(G\mathcal{S}_{1})\ominus(G\mathcal{S}_{2})$.
@@
$\left[\begin{array}[]{c}G_{C}\\ A_{C}\end{array}\right]\Gamma=\left[\begin{array}[]{c}I_{n}\\ 0_{M_{C}\times n}\end{array}\right],$ (25)
@@
with $c_{M^{-}}\triangleq c_{C}-c_{S}$ (see (3)). By Lemma 1 and [23, Thm. 2.1.iv],
@@
$\mathcal{C}_{L}$ =$=\{\xi\ |\ \forall v\in\mathcal{V}_{0},\ \xi+v\in\mathcal{B}_{\infty}(A_{C},b_{C})\}$ (29)
=$=\left\{\xi\ \middle|\ \forall v\in\mathcal{V}_{0},\ A_{C}(\xi+v)=b_{C},\ {\|\xi+v\|}_{\infty}\leq 1\right\}$ (30)
=$=\left\{\xi\ \middle|\ A_{C}\xi=b_{C},\ \forall v\in\mathcal{V}_{0},\ {\|\xi+v\|}_{\infty}\leq 1\right\}$ (31)
=$=\{\xi\ |\ A_{C}\xi=b_{C}\}\cap\left(\{\xi\ |\ {\|\xi\|}_{\infty}\leq 1\}\ominus\mathcal{V}_{0}\right)$ (32)
=$=\left\{\xi\ \middle|A_{C}\xi=b_{C},\ \forall i\in\mathbb{N}_{[1:N_{C}]},\ \forall\delta\in\{-1,1\},\ \delta e_{i}^{\top}\xi\leq 1-\rho_{\mathcal{V}_{0}}\left(\delta e_{i}\right)\right\}$ (33)
=$=\left\{\xi\ \middle|A_{C}\xi=b_{C},\ \forall i\in\mathbb{N}_{[1:N_{C}]},\ |e_{i}^{\top}\xi|\leq 1-\rho_{\mathcal{S}_{0}}\left(\Gamma^{\top}e_{i}\right)\right\}$ (34)
=$=\left\{\xi\ \middle|A_{C}\xi=b_{C},\ \forall i\in\mathbb{N}_{[1:N_{C}]},\ \left|e_{i}^{\top}\xi\right|\leq D_{ii}\right\}$ (35)
Here, (29) follows from (6d) and the definition of $\mathcal{C}_{L}$, (30) follows from (3), (31) follows from the choice of $\mathcal{V}_{0}$, (32) follows from (6d), (33) follows from (8), (34) follows from (4) and from $\rho_{\mathcal{V}_{0}}(\nu)=\rho_{\mathcal{S}_{0}}(\Gamma^{\top}\nu)$ for all $\nu\in\mathbb{R}^{N_{C}}$ [22, Prop. 2], and from symmetry of $\mathcal{S}_{0}$ about the origin implying $\rho_{\mathcal{S}_{0}}(-\mu)=\rho_{\mathcal{S}_{0}}(\mu)$ for all $\mu\in\mathbb{R}^{n}$ by (4), and (35) follows from the definition of $D_{ii}$ (26).
@@
$\mathcal{C}_{L}$ =$=\left\{D\xi\middle|A_{C}D\xi=b_{C},\ {\|\xi\|}_{\infty}\leq 1\right\}=D\mathcal{B}_{\infty}(A_{C}D,b_{C}).$
@@
Theorem 1 inner-approximates the Pontryagin difference (6d) using the support function of the subtrahend $\mathcal{S}$, without requiring the minuend $\mathcal{C}$ to be full-dimensional (Assumption 1). The condition $D_{ii}\geq 0$ for all $i\in\mathbb{N}_{[1:N_{C}]}$ in Theorem 1 may be viewed as requiring $\Gamma$ to additionally satisfy $\Gamma\mathcal{S}_{0}\subseteq\{\xi\ |\|\xi\|_{\infty}\leq 1\}$. Such a choice ensures $\rho_{\mathcal{S}_{0}}\left(\Gamma^{\top}e_{i}\right)\leq 1$, and thereby, $D_{ii}\geq 0$ [21, Ex. 3.35(d)].
We now propose a closed-form expression for $\Gamma$ for full-dimensional minuend $\mathcal{C}$ to address Problem 1.
Theorem 2.
Let Assumption 1 hold for the minuend $\mathcal{C}$. Then, for any convex and compact subtrahend $\mathcal{S}\subset\mathbb{R}^{n}$ that is symmetric about any $c_{S}\in\mathbb{R}^{n}$, the set $\mathcal{M}^{-}$ defined in Theorem 1 with $\Gamma=[G_{C};A_{C}]^{\dagger}[I_{n};0_{M_{C}\times n}]$ inner-approximates $\mathcal{C}\ominus\mathcal{S}$.
@@
Under Assumption 1, $[G_{C};A_{C}]$ has full row rank by Proposition 1. Choose $\Gamma=[G_{C};A_{C}]^{\dagger}[I_{n};0_{M_{C}\times n}]$, and observe that $\Gamma$ satisfies (25), which in turn satisfies the requirements of Theorem 1. ∎
Next, we specialize Theorem 2 for some specific types of symmetric sets, as described in Corollary 1 and Table 1.
@@
Assume the setting of Theorem 2. Then, $D$ in (26) has closed-form expressions for $D_{ii}$ for each $i\in\mathbb{N}_{[1:N_{C}]}$:
1) when $\mathcal{S}$ is a convex union of symmetric intervals (1c),
$D_{ii}=1-{\left\|{e_{i}^{\top}[G_{C};A_{C}]^{\dagger}[G_{S};0_{M_{C}\times N_{S}}]}\right\|}_{\infty},$ (36)
2) when $\mathcal{S}$ is a zonotope (1d),
@@
3) when $\mathcal{S}$ is an ellipsoid (1e),
$D_{ii}=1-{\left\|{e_{i}^{\top}[G_{C};A_{C}]^{\dagger}[G_{S};0_{M_{C}\times N_{S}}]}\right\|}_{2}.$ (38)
@@
Follows from (2.1) and Theorem 2. ∎
Corollary 1 lists some broad classes of subtrahends that admits closed-form expressions for the inner-approximation $\mathcal{M}^{-}$ of the Pontryagin difference $\mathcal{C}\ominus\mathcal{S}$.

### v2

@@ 2.3 Robust controllable set
$\mathcal{K}=\left\{x_{0}\in\mathcal{X}_{0}\ \middle|\ \begin{array}[]{c}\forall t\in\mathbb{N}_{[0:T-1]},\ \exists u_{t}\in\mathcal{U}_{t},\ \forall w_{t}\in\mathcal{W}_{t},\\ x_{t+1}=A_{t}x_{t}+B_{t}u_{t}+F_{t}w_{t},\\ x_{t}\in\mathcal{X}_{t},\ x_{T}\in\mathcal{G}\end{array}\right\}.$ (20)
Informally, the $T$-step RC set $\mathcal{K}\subset\mathbb{R}^{n}$ is the set of initial states that, despite the additive disturbance $w_{t}\in\mathcal{W}_{t}$, can be steered using $u_{t}\in\mathcal{U}_{t}$ to reach the goal set $\mathcal{G}$ at time step $T$, while staying within the state constraints $\mathcal{X}_{t}$ at all intermediate time steps $t\in\mathbb{N}_{[0:T-1]}$. The $T$-step RC set is also known as robust reachability of target tube [4], backward reachable set [9], or backward reach-avoid set [10]. The following set recursion yields $\mathcal{K}=\mathcal{K}_{0}$,
@@
$\operatorname{Pre}(\mathcal{K}_{t+1})$$\triangleq\{x\mid A_{t}x\in(\mathcal{K}_{t+1}\ominus F_{t}\mathcal{W}_{t})\oplus(-B_{t}\mathcal{U}_{t})\},$ (21b)
with $\mathcal{K}_{T}\triangleq\mathcal{G}$. Observe that (21) uses all set operations in (6). For H-Rep/V-Rep polytopes $\mathcal{X},\mathcal{G},\mathcal{U}$, and $\mathcal{W}$, we can compute RC sets using existing results [24, 10, 3]. However, when using H-Rep/V-Rep polytopes for high-dimensional systems (16) or over long horizons $T$, we face numerical challenges when implementing (21), since it requires a combination of Minkowski sum and intersection operations. See [3, 16, 17, 18] for a detailed discussion.
On the other hand, constrained zonotopes provide a tractable, exact solution when $\mathcal{W}=\emptyset$ using (7[5, 6, 9]. Also, for a zonotopic $\mathcal{W}$[9] uses the two-stage approach reviewed in Sec. 2.2 to propose an inner-approximation of $\mathcal{K}$. In this paper, we use our solution to Prob. 1 to inner-approximate $\mathcal{K}$ for a broader class of sets $\mathcal{W}$.
@@
Compute a constrained zonotope $\mathcal{K}^{-}$ that inner-approximates $\mathcal{K}$ defined in Defn. 2, where for every $t\in\mathbb{N}_{[0:T-1]}$, $\mathcal{U}_{t}$ and $\mathcal{G}$ are polytopes, $\mathcal{W}_{t}$ are symmetric, convex, and compact sets, and either 1) $\mathcal{X}_{t}$ are convex polyhedra and $A_{t}$ are invertible, or 2) $\mathcal{X}_{t}$ are polytopes. Additionally, characterize the representation complexity of $\mathcal{K}^{-}$ in both cases.
In both settings considered, the exact RC sets are known to be polytopes [3], and hence are representable via constrained zonotopes. We also use the proposed outer-approximation to the Pontryagin difference to outer-approximate $\mathcal{K}_{t}$ when using (21).
@@ 3 Inner-approximation of Pontryagin difference
Given a constrained zonotopic minuend $\mathcal{C}\subset\mathbb{R}^{n}$ and a symmetric, convex, and compact subtrahend $\mathcal{S}\subset\mathbb{R}^{n}$, we characterize a constrained zonotope $\mathcal{M}^{-}\subset\mathbb{R}^{n}$ where
@@
$G_{M^{-}}$ $=G_{C}D,$ $c_{M^{-}}$ $=c_{C}-c_{S},$ (23a)
$A_{M^{-}}$ $=A_{C}D,$ $b_{M^{-}}$ $=b_{C},$ (23b)
for a specific $c_{S}\in\mathcal{S}$, similarly to (14). The representation complexity of $\mathcal{M}^{-}$ is $\mathscr{C}(\mathcal{M}^{-})=({M_{C}},\mathscr{O}_{{C}})=\mathscr{C}(\mathcal{C})$.
In this section, we propose two constrained zonotope representations that allow us to address Prob. 1, and relate our approach with the two-stage approach in [9] for zonotopic subtrahends. Table 1 summarizes the computation of $D$ for various types of sets.
Subtrahend $\mathcal{S}$ Definition of the diagonal matrix $D=\operatorname{diag}_{i\in\mathbb{N}_{[1:{N_{C}}]}}(D_{ii})$ Result Convex and compact $\mathcal{S}$, symmetric about $c_{S}\in\mathbb{R}^{n}$ $D_{ii}=1-\rho_{\mathcal{S}_{0}}\left({e_{i}^{\top}{[G_{C};A_{C}]^{\dagger}[I_{n};0_{{M_{C}}\times n}]}}\right)$ for $\mathcal{S}_{0}\triangleq\mathcal{S}-c_{S}$ Prop. 2  Zonotope $\mathcal{S}$ (3a) $D_{ii}=1-{\left\|{e_{i}^{\top}[G_{C};A_{C}]^{\dagger}[G_{S};0_{{M_{C}}\times{N_{S}}}]}\right\|}_{1}$ Ellipsoid $\mathcal{S}$ (3b) $D_{ii}=1-{\left\|{e_{i}^{\top}[G_{C};A_{C}]^{\dagger}[G_{S};0_{{M_{C}}\times n}]}\right\|}_{2}$ Corr. 1 Convex union of symmetric intervals $\mathcal{S}$ (3c) $D_{ii}=1-{\left\|{e_{i}^{\top}[G_{C};A_{C}]^{\dagger}[G_{S};0_{{M_{C}}\times{N_{S}}}]}\right\|}_{\infty}$
@@
3.1 Full-dimensional constrained zonotopes
Figure 1: Relationship between various representations discussed in the paper. For any pair of representations $R_{1},R_{2}$, a dashed arrow from $R_{1}$ to $R_{2}$ shows that $R_{1}$ is also $R_{2}$.
We study two constrained zonotope representations that are guaranteed to exist for full-dimensional polytopes.
@@ 3.1 Full-dimensional constrained zonotopes
(MinRow Representation) A constrained zonotope $\mathcal{C}=(G_{C},c_{C},A_{C},b_{C})$ is a MinRow representation when $[G_{C};A_{C}]$ has full row rank.
Definition 4.
(Invertible representation) A MinRow representation $\mathcal{C}$ is an Invertible representation when $n+{M_{C}}={N_{C}}$, i.e., $\mathscr{O}_{{\mathcal{C}}}=1$.
By definitions, the matrix $[G_{C};A_{C}]\in\mathbb{R}^{(n+{M_{C}})\times{N_{C}}}$ has a well-defined right pseudoinverse ${[G_{C};A_{C}]}^{\dagger}$ for a MinRow representation, and $[G_{C};A_{C}]$ is invertible for an Invertible representation. Fig. 1 illustrates the relationships between the representations based on Prop. 1.
Algorithm 1 Compute MinRow representation
1:Constrained zonotope $(G_{C},c_{C},A_{C},b_{C})$ that represents a full-dimensional polytope
2:MinRow representation $(G_{C}^{\prime},c_{C}^{\prime},A_{C}^{\prime},b_{C}^{\prime})$
3:Set $G_{C}^{\prime}\leftarrow G_{C}$ and $c_{C}^{\prime}\leftarrow c_{C}$.
4:Get $A_{C}^{\prime},b_{C}^{\prime}$ from rows of $[A_{C},b_{C}]$ where $[A_{C}^{\prime},b_{C}^{\prime}]$ has full row rank and $\operatorname{rank}([A_{C}^{\prime},b_{C}^{\prime}])=\operatorname{rank}([A_{C},b_{C}])$.
Algorithm 2 Compute Invertible representation
1:H-Rep polytope $\mathcal{C}=\{x\mid H_{C}x\leq k_{C}\}\subset\mathbb{R}^{n}$ that represents a full-dimensional polytope
2:Invertible representation $(G_{C},c_{C},A_{C},b_{C})$
3:Define a zonotope $\mathcal{Z}=(G_{Z},c_{Z})$ (3a) with $G_{Z}\triangleq{\operatorname{diag}\left(\frac{u-l}{2}\right)\in\mathbb{R}^{n\times n},c_{Z}\triangleq\frac{u+l}{2}}\in\mathbb{R}^{n}$ for $l,u\in\mathbb{R}^{n}$,
$l_{i}=-\rho_{\mathcal{C}}(-e_{i}),\ \ u_{i}=\rho_{\mathcal{C}}(e_{i}),\quad\forall i\in\mathbb{N}_{[1:n]}.$ (24)
4:Define $\sigma\in\mathbb{R}^{{L_{C}}}$ with $\sigma_{i}=-\rho_{\mathcal{Z}}(-h_{i})$ for each $i\in\mathbb{N}_{[1:{L_{C}}]}$ and $H_{C}=[h_{1}^{\top};h_{2}^{\top};\ldots;h_{{L_{C}}}^{\top}]$.
5:Compute the Invertible representation for $\mathcal{C}$ with
$G_{C}$ $=[G_{Z},0_{n\times{L_{C}}}],$ $c_{C}$ $=c_{Z},$
$A_{C}$ $=\left[{H_{C}G_{Z},\operatorname{diag}\left(\frac{\sigma-k}{2}\right)}\right],$ $b_{C}$ $=\frac{\sigma+k}{2}-H_{C}c_{Z}.$
@@
The following statements hold:
1) Every representation $(G_{C},c_{C},A_{C},b_{C})$ of a full-dimensional polytope $\mathcal{C}$ satisfies $\operatorname{rank}(G_{C})=n\leq{N_{C}}$.
2) A polytope is full-dimensional if and only if the polytope is non-empty and it has a MinRow representation.
3) Algo. 1 generates a MinRow representation from any full-dimensional constrained zonotope (1b).
4) Algo. 2 generates an Invertible representation from any full-dimensional H-Rep polytope (1a).
5) Given an Invertible representation $(G_{C},c_{C},A_{C},b_{C})$, the corresponding H-Rep polytope (1a) is given by,
$H_{C}$ $=\left[I_{{N_{C}}};-I_{{N_{C}}}\right]{\left[G_{C};A_{C}\right]}^{-1}\left[I_{n};0_{{M_{C}}\times n}\right],$ (25a)
$k_{C}$ $=1_{2{N_{C}}\times 1}-\left[I_{{N_{C}}};-I_{{N_{C}}}\right]{\left[G_{C};A_{C}\right]}^{-1}\left[-c;b\right].$ (25b)
See Sec. A.1 for the proof of Prop. 1. Step 4 in Algo. 1 removes redundant equalities in $\{\xi\mid A_{C}\xi=b_{C}\}$. See mpt_minAffineRep in MPT3 [24] for an implementation.
By Prop. 1.4, an Invertible representation also exists for every full-dimensional constrained zonotope. However, it is unclear if we can efficiently compute such a representation from an arbitrary full-dimensional constrained zonotope, as done in Algo. 1.
Remark 1.
The rest of the paper assumes the use of a MinRow representation for any full-dimensional constrained zonotope, obtained using either Algo. 1 or 2.
@@ 3.2 Inner-approximation of the Pontryagin difference
Thm. 1 tackles the inner-approximation part of Prob. 1.
@@
$[G_{C};A_{C}]\Gamma=[I_{n};0_{{M_{C}}\times n}],$ (26)
@@
with $c_{M^{-}}$ in (23). Recall that for any set $\mathcal{A},\mathcal{B}\subset\mathbb{R}^{n}$ and a matrix $M\in\mathbb{R}^{m\times n}$ with $m<n$, $M(\mathcal{A}\ominus\mathcal{B})\subseteq(M\mathcal{A})\ominus(M\mathcal{B})$ [9, Lem. 2(ii)]. From [27, Thm. 2.1.iv],
@@
$\mathcal{C}_{L}$ =$=\{\xi\mid\forall v\in\mathcal{V}_{0},\ \xi+v\in\mathcal{B}_{\infty}(A_{C},b_{C})\}$ (30)
=$=\{\xi\mid\forall v\in\mathcal{V}_{0},\ A_{C}(\xi+v)=b_{C},\ {\|\xi+v\|}_{\infty}\leq 1\}$ (31)
=$=\{\xi\mid A_{C}\xi=b_{C},\ \forall v\in\mathcal{V}_{0},\ {\|\xi+v\|}_{\infty}\leq 1\}$ (32)
=$=\{\xi\mid A_{C}\xi=b_{C}\}\cap\left(\{\xi\mid{\|\xi\|}_{\infty}\leq 1\}\ominus\mathcal{V}_{0}\right)$ (33)
=$=\left\{\xi\ \middle|\begin{array}[]{c}A_{C}\xi=b_{C},\ \forall i\in\mathbb{N}_{[1:{N_{C}}]},\ \forall\delta\in\{-1,1\},\\ \delta e_{i}^{\top}\xi\leq 1-\rho_{\mathcal{V}_{0}}\left(\delta e_{i}\right)\end{array}\right\}$ (36)
=$=\left\{\xi\ \middle|\ \begin{array}[]{c}A_{C}\xi=b_{C},\ \forall i\in\mathbb{N}_{[1:{N_{C}}]},\\ |e_{i}^{\top}\xi|\leq 1-\rho_{\mathcal{S}_{0}}\left(\Gamma^{\top}e_{i}\right)\end{array}\right\}$ (39)
=$=\left\{\xi\ \middle|A_{C}\xi=b_{C},\ \forall i\in\mathbb{N}_{[1:{N_{C}}]},\ \left|e_{i}^{\top}\xi\right|\leq D_{ii}\right\}.$ (40)
Here, (30) follows from (6d) and the definition of $\mathcal{C}_{L}$, (31) follows from (2), (32) follows from the choice of $\mathcal{V}_{0}$, (33) follows from (6d), (36) follows from (8), (39) follows from (4) and from $\rho_{\mathcal{V}_{0}}(\nu)=\rho_{\mathcal{S}_{0}}(\Gamma^{\top}\nu)$ for all $\nu\in\mathbb{R}^{{N_{C}}}$ [26, Prop. 2], and from symmetry of $\mathcal{S}_{0}$ about the origin implying $\rho_{\mathcal{S}_{0}}(-\mu)=\rho_{\mathcal{S}_{0}}(\mu)$ for all $\mu\in\mathbb{R}^{n}$ by (4), and (40) follows from the definition of $D_{ii}$ (27).
@@
$\mathcal{C}_{L}$ =$=\{D\xi\mid A_{C}D\xi=b_{C},\ {\|\xi\|}_{\infty}\leq 1\}=D\mathcal{B}_{\infty}(A_{C}D,b_{C}).$
@@
Thm. 1 inner-approximates the Pontryagin difference (6d) using the support function of the subtrahend $\mathcal{S}$. The condition $D_{ii}\geq 0$ for all $i\in\mathbb{N}_{[1:{N_{C}}]}$ in Thm. 1 may be viewed as requiring $\Gamma$ to additionally satisfy $\Gamma\mathcal{S}_{0}\subseteq\{\xi\mid\|\xi\|_{\infty}\leq 1\}$. Such a choice ensures $\rho_{\mathcal{S}_{0}}\left(\Gamma^{\top}e_{i}\right)\leq 1$, and thereby, $D_{ii}\geq 0$ [22, Ex. 3.35(d)].
Proposition 2.
For a full-dimensional, constrained zonotopic minuend $\mathcal{C}$ and a convex and compact subtrahend $\mathcal{S}\subset\mathbb{R}^{n}$ that is symmetric about any $c_{S}\in\mathbb{R}^{n}$, the constrained zonotope $\mathcal{M}^{-}$ defined in Thm. 1 with $\Gamma=[G_{C};A_{C}]^{\dagger}[I_{n};0_{{M_{C}}\times n}]$ inner-approximates $\mathcal{C}\ominus\mathcal{S}$.
@@
Follows from Rem. 1 and Thm. 1. ∎
Prop. 2 addresses Prob. 1. Additionally, we use the following two observations to illustrate that the assumption of full-dimensionality in Prop. 2 is not restrictive. First, $\mathcal{C}\ominus\mathcal{S}=\emptyset$, whenever the affine dimension of $\mathcal{S}$ is strictly greater than the affine dimension of $\mathcal{C}$ by (6d), since $x+\mathcal{S}$ can not be contained in $\mathcal{C}$ for any $x\in\mathbb{R}^{n}$ in such a scenario. Second, consider the case where both minuend and subtrahend are not full-dimensional and there exists a matrix $T\in\mathbb{R}^{n\times n^{\prime}}$ with $n>n^{\prime}$ and $\text{rank}(T)=n^{\prime}$ such that $\mathcal{C}=T\mathcal{C}^{\prime}$ and $\mathcal{S}=T\mathcal{S}^{\prime}$ for any $\mathcal{C}^{\prime},\mathcal{S}^{\prime}\subset\mathbb{R}^{n^{\prime}}$. Then, by [27, Thm. 2.1.viii],
In such a case, it suffices to assume that $\mathcal{C}^{\prime}$ is full-dimensional in order to obtain a constrained zonotope inner-approximation of $\mathcal{C}\ominus\mathcal{S}$.
@@
For a full-dimensional constrained zonotopic minuend $\mathcal{C}$, $D$ in (27) has closed-form expressions for $D_{ii}$ for each $i\in\mathbb{N}_{[1:{N_{C}}]}$:
1) when $\mathcal{S}$ is a zonotope (3a),
@@
2) when $\mathcal{S}$ is an ellipsoid (3b),
$D_{ii}=1-{\left\|{e_{i}^{\top}[G_{C};A_{C}]^{\dagger}[G_{S};0_{{M_{C}}\times{N_{S}}}]}\right\|}_{2},$ (43)
3) when $\mathcal{S}$ is a convex union of symmetric intervals (3c),
$D_{ii}=1-{\left\|{e_{i}^{\top}[G_{C};A_{C}]^{\dagger}[G_{S};0_{{M_{C}}\times{N_{S}}}]}\right\|}_{\infty}.$ (44)
@@
Follows from (5) and Prop. 2. ∎
Corr. 1 lists some broad classes of subtrahends that admits closed-form expressions for the inner-approximation $\mathcal{M}^{-}$ of the Pontryagin difference $\mathcal{C}\ominus\mathcal{S}$.